# FeSi Surface Ferromagnetism Interplay between 2D ferromagnetism and transport at the surface of FeSi

### v1

@@ Acknowledgements.
- Zhu et al. [2013] Z.-H. Zhu, A. Nicolaou, G. Levy, N. P. Butch, P. Syers, X. F. Wang, J. Paglione, G. A. Sawatzky, I. S. Elfimov, and A. Damascelli, Phys. Rev. Lett. 111, 216402 (2013).
@@
- Eo et al. [2023] Y. S. Eo, K. Avers, J. A. Horn, H. Yoon, S. R. Saha, A. Suarez, M. S. Fuhrer, and J. Paglione, Applied Physics Letters 122, 233102 (2023).
- Sarma and Hwang [2014] S. D. Sarma and E. H. Hwang, Phys. Rev. B 89, 235423 (2014).
- Zabrodskii [2001] A. G. Zabrodskii, Philosophical Magazine B 81, 1131 (2001), https://doi.org/10.1080/13642810108205796 .
@@
- Sarma and Hwang [2005] S. D. Sarma and E. Hwang, Solid State Communications 135, 579 (2005), fundamental Optical and Quantum Effects in Condensed Matter.
@@
- Huang et al. [2022a] Y. Huang, B. Skinner, and B. I. Shklovskii, Journal of Experimental and Theoretical Physics 135, 409 (2022a).
- Huang et al. [2022b] Y. Huang, Y. He, B. Skinner, and B. I. Shklovskii, Phys. Rev. B 105, 054206 (2022b).
@@
- Schlesinger et al. [1993] Z. Schlesinger, Z. Fisk, H.-T. Zhang, M. B. Maple, J. DiTusa, and G. Aeppli, Phys. Rev. Lett. 71, 1748 (1993).
@@
- Buschinger et al. [1997] B. Buschinger, C. Geibel, F. Steglich, D. Mandrus, D. Young, J. Sarrao, and Z. Fisk, Physica B: Condensed Matter 230-232, 784 (1997), proceedings of the International Conference on Strongly Correlated Electron Systems.
@@
- Degiorgi et al. [1995a] L. Degiorgi, M. Hunt, H. R. Ott, and Z. Fisk, PHYSICA B-CONDENSED MATTER 206, 810 (1995a).
@@
- Glushkov et al. [2000] V. Glushkov, N. Sluchanko, S. Demishev, M. Kondrin, A. Pronin, K. Petukhov, Y. Bruynseraede, V. Moshchalkov, and A. Menovsky, PHYSICA B 284, 1179 (2000), 22nd International Conference on Low Temperature Physics, HELSINKI UNIV TECHNOL, HELSINKI, FINLAND, AUG 04-11, 1999.
@@
- Lisunov et al. [1996] K. Lisunov, E. Arushanov, C. Kloc, J. Broto, J. Leotin, H. Rokoto, M. Respaud, and E. Bucher, PHYSICA B 229, 37 (1996).
@@
- Mihalik et al. [1996] M. Mihalik, M. Timko, P. Samuely, N. Tomašovičova-Hudákova, P. Szabó, and A. Menovsky, Journal of Magnetism and Magnetic Materials 157-158, 637 (1996), european Magnetic Materials and Applications Conference.
@@
- Paschen et al. [1997] S. Paschen, E. Felder, M. A. Chernikov, L. Degiorgi, H. Schwer, H. R. Ott, D. P. Young, J. L. Sarrao, and Z. Fisk, Phys. Rev. B 56, 12916 (1997).
@@
- Changdar et al. [2020] S. Changdar, S. Aswartham, A. Bose, Y. Kushnirenko, G. Shipunov, N. C. Plumb, M. Shi, A. Narayan, B. Büchner, and S. Thirupathaiah, Phys. Rev. B 101, 235105 (2020).
@@
- Lunkenheimer et al. [1995] P. Lunkenheimer, G. Knebel, R. Viana, and A. Loidl, Solid State Communications 93, 891 (1995).
@@
- Degiorgi et al. [1995b] L. Degiorgi, M. Hunt, H. Ott, and Z. Fisk, Physica B: Condensed Matter 206-207, 810 (1995b), proceedings of the International Conference on Strongly Correlated Electron Systems.
@@
- Ou-Yang et al. [2017] T. Ou-Yang, Y. Zhuang, B. Ramachandran, W. Chen, G. Shu, C. Hu, F. Chou, and Y. Kuo, Journal of Alloys and Compounds 702, 92 (2017).
@@
- Delaire et al. [2015] O. Delaire, I. I. Al-Qasir, A. F. May, C. W. Li, B. C. Sales, J. L. Niedziela, J. Ma, M. Matsuda, D. L. Abernathy, and T. Berlijn, Phys. Rev. B 91, 094307 (2015).
- Mermin and Wagner [1966] N. D. Mermin and H. Wagner, Phys. Rev. Lett. 17, 1133 (1966).
- Gong et al. [2017] C. Gong, L. Lin, L. Zhenglu, J. Huiwen, A. Stern, X. Yang, C. Ting, B. Wei, C. Wang, Y. Wang, Z. Q. Qui, R. J. Cava, S. G. Louie, X. Jing, and Z. Xiang, Nature 546, 265 (2017).
- Koyama et al. [2000] K. Koyama, T. Goto, T. Kanomata, R. Note, and Y. Takahashi, Journal of the Physical Society of Japan 69, 219 (2000).
- Nagaosa et al. [2010] N. Nagaosa, J. Sinova, S. Onoda, A. H. MacDonald, and N. P. Ong, Rev. Mod. Phys. 82, 1539 (2010).
- Kim et al. [2013] D. J. Kim, S. Thomas, T. Grant, J. Botimer, Z. Fisk, and J. Xia, Sci. Rep. 3, 3150 (2013).
- Balasubramanian et al. [2020] B. Balasubramanian, P. Manchanda, R. Pahari, Z. Chen, W. Zhang, S. R. Valloppilly, X. Li, A. Sarella, L. Yue, A. Ullah, P. Dev, D. A. Muller, R. Skomski, G. C. Hadjipanayis, and D. J. Sellmyer, Phys. Rev. Lett. 124, 057201 (2020).
- Münzer et al. [2010] W. Münzer, A. Neubauer, T. Adams, S. Mühlbauer, C. Franz, F. Jonietz, R. Georgii, P. Böni, B. Pedersen, M. Schmidt, A. Rosch, and C. Pfleiderer, Phys. Rev. B 81, 041203 (2010).
- Bauer et al. [2016] A. Bauer, M. Garst, and C. Pfleiderer, Phys. Rev. B 93, 235144 (2016).
- Ohtsuka et al. [2021] Y. Ohtsuka, N. Kanazawa, M. Hirayama, A. Matsui, T. Nomoto, R. Arita, T. Nakajima, T. Hanashima, V. Ukleev, H. Aoki, M. Mogi, K. Fujiwara, A. Tsukazaki, M. Ichikawa, M. Kawasaki, and Y. Tokura, Science Advances 7, eabj0498 (2021).
- Manyala et al. [2009] N. Manyala, B. D. Ngom, A. C. Beye, R. Bucher, M. Maaza, A. Strydom, A. Forbes, J. Johnson, A. T. Charlie, and J. F. DiTusa, Applied Physics Letters 94, 232503 (2009).
@@
- Ángel R Ruiz et al. [2018] Ángel R Ruiz, J. Hernández-Pérez, L. F. Fonseca, M. José-Yacamán, E. Ortega, and A. Ponce, Nanotechnology 30, 014001 (2018).
@@
- Manyala et al. [2004] N. Manyala, Y. Sidis, J. F. DiTusa, G. Aeppli, D. P. Young, and Z. Fisk, Nature Materials 3, 255 (2004).
@@
- Breindel et al. [2023] A. Breindel, Y. Deng, C. M. Moir, Y. Fang, S. Ran, H. Lou, S. Li, Q. Zeng, L. Shu, C. T. Wolowiec, I. K. Schuller, P. F. S. Rosa, Z. Fisk, J. Singleton, and M. B. Maple, Proceedings of the National Academy of Sciences 120, e2216367120 (2023).
@@
- Deng et al. [2023] Y. Deng, Y. Yan, H. Wang, E. Lee-Wong, C. M. Moir, Y. Fang, W. Xie, and M. B. Maple, Phys. Rev. B 108, 115158 (2023).
@@
- Brando et al. [2016] M. Brando, D. Belitz, F. M. Grosche, and T. R. Kirkpatrick, Rev. Mod. Phys. 88, 025006 (2016).
- Saha et al. [2022] S. Saha, A. Dutta, S. Gupta, S. Bandyopadhyay, and I. Das, Phys. Rev. B 105, 214407 (2022).
- Das et al. [2018] K. Das, N. Banu, I. Das, and B. Dev, Solid State Communications 274, 36 (2018).
@@
- Ohnuma et al. [2012] I. Ohnuma, S. Abe, S. Shimenouchi, T. Omori, R. Kainuma, and K. Ishida, ISIJ International 52, 540 (2012).
@@
- Alexandrov et al. [2015] V. Alexandrov, P. Coleman, and O. Erten, Phys. Rev. Lett. 114, 177202 (2015).
@@
- Nakajima et al. [2016] Y. Nakajima, P. Syers, X. Wang, R. Wang, and J. Paglione, Nature Physics 12, 213 (2016).
@@
- Sun et al. [2017] Y. Sun, Z. Zhuo, X. Wu, and J. Yang, Nano Letters 17, 2771 (2017), pMID: 28441496, https://doi.org/10.1021/acs.nanolett.6b04884 .

### v2

@@ Acknowledgements.
- Zhu et al. [2013] Z.-H. Zhu, A. Nicolaou, G. Levy, N. P. Butch, P. Syers, X. F. Wang, J. Paglione, G. A. Sawatzky, I. S. Elfimov, and A. Damascelli, Polarity-Driven Surface Metallicity in , Phys. Rev. Lett. 111, 216402 (2013).
@@
- Eo et al. [2023] Y. S. Eo, K. Avers, J. A. Horn, H. Yoon, S. R. Saha, A. Suarez, M. S. Fuhrer, and J. Paglione, Extraordinary bulk-insulating behavior in the strongly correlated materials FeSi and FeSb2, Applied Physics Letters 122, 233102 (2023).
- Sarma and Hwang [2014] S. D. Sarma and E. H. Hwang, Two-dimensional metal-insulator transition as a strong localization induced crossover phenomenon, Phys. Rev. B 89, 235423 (2014).
- Zabrodskii [2001] A. G. Zabrodskii, The coulomb gap: The view of an experimenter, Philosophical Magazine B 81, 1131 (2001), https://doi.org/10.1080/13642810108205796 .
@@
- Sarma and Hwang [2005] S. D. Sarma and E. Hwang, The so-called two dimensional metal–insulator transition, Solid State Communications 135, 579 (2005), fundamental Optical and Quantum Effects in Condensed Matter.
@@
- Huang et al. [2022a] Y. Huang, B. Skinner, and B. I. Shklovskii, Conductivity of Two-Dimensional Small Gap Semiconductors and Topological Insulators in Strong Coulomb Disorder, Journal of Experimental and Theoretical Physics 135, 409 (2022a).
- Huang et al. [2022b] Y. Huang, Y. He, B. Skinner, and B. I. Shklovskii, Conductivity of two-dimensional narrow gap semiconductors subjected to strong Coulomb disorder, Phys. Rev. B 105, 054206 (2022b).
@@
- Schlesinger et al. [1993] Z. Schlesinger, Z. Fisk, H.-T. Zhang, M. B. Maple, J. DiTusa, and G. Aeppli, Unconventional charge gap formation in FeSi, Phys. Rev. Lett. 71, 1748 (1993).
@@
- Buschinger et al. [1997] B. Buschinger, C. Geibel, F. Steglich, D. Mandrus, D. Young, J. Sarrao, and Z. Fisk, Transport properties of fesi, Physica B: Condensed Matter 230-232, 784 (1997), proceedings of the International Conference on Strongly Correlated Electron Systems.
@@
- Degiorgi et al. [1995a] L. Degiorgi, M. Hunt, H. R. Ott, and Z. Fisk, Transport and optical properties of FeSi, PHYSICA B-CONDENSED MATTER 206, 810 (1995a).
@@
- Glushkov et al. [2000] V. Glushkov, N. Sluchanko, S. Demishev, M. Kondrin, A. Pronin, K. Petukhov, Y. Bruynseraede, V. Moshchalkov, and A. Menovsky, Low-temperature transport anomalies in FeSi, PHYSICA B 284, 1179 (2000), 22nd International Conference on Low Temperature Physics, HELSINKI UNIV TECHNOL, HELSINKI, FINLAND, AUG 04-11, 1999.
@@
- Lisunov et al. [1996] K. Lisunov, E. Arushanov, C. Kloc, J. Broto, J. Leotin, H. Rokoto, M. Respaud, and E. Bucher, Conductivity and magnetoresistance of FeSi in the Anderson-localized regime, PHYSICA B 229, 37 (1996).
@@
- Mihalik et al. [1996] M. Mihalik, M. Timko, P. Samuely, N. Tomašovičova-Hudákova, P. Szabó, and A. Menovsky, Magnetic properties and gap formation in FeSi, Journal of Magnetism and Magnetic Materials 157-158, 637 (1996), european Magnetic Materials and Applications Conference.
@@
- Paschen et al. [1997] S. Paschen, E. Felder, M. A. Chernikov, L. Degiorgi, H. Schwer, H. R. Ott, D. P. Young, J. L. Sarrao, and Z. Fisk, Low-temperature transport, thermodynamic, and optical properties of FeSi, Phys. Rev. B 56, 12916 (1997).
@@
- Changdar et al. [2020] S. Changdar, S. Aswartham, A. Bose, Y. Kushnirenko, G. Shipunov, N. C. Plumb, M. Shi, A. Narayan, B. Büchner, and S. Thirupathaiah, Electronic structure studies of FeSi: A chiral topological system, Phys. Rev. B 101, 235105 (2020).
@@
- Lunkenheimer et al. [1995] P. Lunkenheimer, G. Knebel, R. Viana, and A. Loidl, Hopping conductivity in FeSi, Solid State Communications 93, 891 (1995).
@@
- Degiorgi et al. [1995b] L. Degiorgi, M. Hunt, H. Ott, and Z. Fisk, Transport and optical properties of FeSi, Physica B: Condensed Matter 206-207, 810 (1995b), proceedings of the International Conference on Strongly Correlated Electron Systems.
@@
- Ou-Yang et al. [2017] T. Ou-Yang, Y. Zhuang, B. Ramachandran, W. Chen, G. Shu, C. Hu, F. Chou, and Y. Kuo, Effect of Co substitution on thermoelectric properties of FeSi, Journal of Alloys and Compounds 702, 92 (2017).
@@
- Delaire et al. [2015] O. Delaire, I. I. Al-Qasir, A. F. May, C. W. Li, B. C. Sales, J. L. Niedziela, J. Ma, M. Matsuda, D. L. Abernathy, and T. Berlijn, Heavy-impurity resonance, hybridization, and phonon spectral functions in , Phys. Rev. B 91, 094307 (2015).
@@
- Ángel R Ruiz et al. [2018] Ángel R Ruiz, J. Hernández-Pérez, L. F. Fonseca, M. José-Yacamán, E. Ortega, and A. Ponce, Single nanowire measurements of room temperature ferromagnetism in FeSi nanowires and the effects of Mn-doping, Nanotechnology 30, 014001 (2018).
- Ohtsuka et al. [2021] Y. Ohtsuka, N. Kanazawa, M. Hirayama, A. Matsui, T. Nomoto, R. Arita, T. Nakajima, T. Hanashima, V. Ukleev, H. Aoki, M. Mogi, K. Fujiwara, A. Tsukazaki, M. Ichikawa, M. Kawasaki, and Y. Tokura, Emergence of spin-orbit coupled ferromagnetic surface state derived from Zak phase in a nonmagnetic insulator FeSi, Science Advances 7, eabj0498 (2021).
- Burmistrov and Repin [2018] I. S. Burmistrov and E. V. Repin, Quantum corrections to conductivity of disordered electrons due to inelastic scattering off magnetic impurities, Phys. Rev. B 98, 045414 (2018).
- Mitra et al. [2007] P. Mitra, R. Misra, A. F. Hebard, K. A. Muttalib, and P. Wölfle, Weak-Localization Correction to the Anomalous Hall Effect in Polycrystalline Fe Films, Phys. Rev. Lett. 99, 046804 (2007).
- Lin et al. [1996] Y. K. Lin, T. R. Novet, D. C. Johnson, and J. M. Valles, Magnetotransport studies of strongly disordered annealed amorphous Fe/Si multilayers, Phys. Rev. B 53, 4796 (1996).
- Mermin and Wagner [1966] N. D. Mermin and H. Wagner, Absence of Ferromagnetism or Antiferromagnetism in One- or Two-Dimensional Isotropic Heisenberg Models, Phys. Rev. Lett. 17, 1133 (1966).
- Gong et al. [2017] C. Gong, L. Lin, L. Zhenglu, J. Huiwen, A. Stern, X. Yang, C. Ting, B. Wei, C. Wang, Y. Wang, Z. Q. Qui, R. J. Cava, S. G. Louie, X. Jing, and Z. Xiang, Discovery of intrinsic ferromagnetism in two-dimensional van der Waals crystals, Nature 546, 265 (2017).
- Koyama et al. [2000] K. Koyama, T. Goto, T. Kanomata, R. Note, and Y. Takahashi, Nonlinear Magnetization Process of Single-Crystalline FeSi, Journal of the Physical Society of Japan 69, 219 (2000).
- Renard et al. [2005] V. T. Renard, I. V. Gornyi, O. A. Tkachenko, V. A. Tkachenko, Z. D. Kvon, E. B. Olshanetsky, A. I. Toropov, and J.-C. Portal, Quantum corrections to the conductivity and Hall coefficient of a two-dimensional electron gas in a dirty quantum well: From the diffusive to the ballistic regime, Phys. Rev. B 72, 075313 (2005).
- Chen et al. [2011] J.-H. Chen, L. Li, W. G. Cullen, E. D. Williams, and M. S. Fuhrer, Tunable Kondo effect in graphene with defects, Nature Physics 7, 535 (2011).
- Nagaosa et al. [2010] N. Nagaosa, J. Sinova, S. Onoda, A. H. MacDonald, and N. P. Ong, Anomalous Hall effect, Rev. Mod. Phys. 82, 1539 (2010).
- Kim et al. [2013] D. J. Kim, S. Thomas, T. Grant, J. Botimer, Z. Fisk, and J. Xia, Surface Hall Effect and Nonlocal Transport in SmB6: Evidence for Surface Conduction, Sci. Rep. 3, 3150 (2013).
- Balasubramanian et al. [2020] B. Balasubramanian, P. Manchanda, R. Pahari, Z. Chen, W. Zhang, S. R. Valloppilly, X. Li, A. Sarella, L. Yue, A. Ullah, P. Dev, D. A. Muller, R. Skomski, G. C. Hadjipanayis, and D. J. Sellmyer, Chiral Magnetism and High-Temperature Skyrmions in B20-Ordered Co-Si, Phys. Rev. Lett. 124, 057201 (2020).
- Münzer et al. [2010] W. Münzer, A. Neubauer, T. Adams, S. Mühlbauer, C. Franz, F. Jonietz, R. Georgii, P. Böni, B. Pedersen, M. Schmidt, A. Rosch, and C. Pfleiderer, Skyrmion lattice in the doped semiconductor , Phys. Rev. B 81, 041203 (2010).
- Bauer et al. [2016] A. Bauer, M. Garst, and C. Pfleiderer, History dependence of the magnetic properties of single-crystal , Phys. Rev. B 93, 235144 (2016).
@@
- Manyala et al. [2004] N. Manyala, Y. Sidis, J. F. DiTusa, G. Aeppli, D. P. Young, and Z. Fisk, Large anomalous Hall effect in a silicon-based magnetic semiconductor, Nature Materials 3, 255 (2004).
- Beille et al. [1981] J. Beille, J. Voiron, F. Towfiq, M. Roth, and Z. Y. Zhang, Helimagnetic structure of the FexCo1-xSi alloys, Journal of Physics F: Metal Physics 11, 2153 (1981).
- Takeda et al. [2009] M. Takeda, Y. Endoh, K. Kakurai, Y. Onose, J. Suzuki, and Y. Tokura, Nematic-to-Smectic Transition of Magnetic Texture in Conical State, Journal of the Physical Society of Japan 78, 093704 (2009), https://doi.org/10.1143/JPSJ.78.093704 .
- Prakash et al. [2007] R. Prakash, R. J. Choudhary, L. S. S. Chandra, N. Lakshmi, and D. M. Phase, Electrical and magnetic transport properties of Fe3O4 thin films on a GaAs(100) substrate, Journal of Physics: Condensed Matter 19, 486212 (2007).
- Hori et al. [2023] T. Hori, N. Kanazawa, M. Hirayama, K. Fujiwara, A. Tsukazaki, M. Ichikawa, M. Kawasaki, and Y. Tokura, A Noble-Metal-Free Spintronic System with Proximity-Enhanced Ferromagnetic Topological Surface State of FeSi above Room Temperature, Advanced Materials 35, 2206801 (2023), https://onlinelibrary.wiley.com/doi/pdf/10.1002/adma.202206801 .
- Hori et al. [2024] T. Hori, N. Kanazawa, K. Matsuura, H. Ishizuka, K. Fujiwara, A. Tsukazaki, M. Ichikawa, M. Kawasaki, F. Kagawa, M. Hirayama, and Y. Tokura, Strongly pinned skyrmionic bubbles and higher-order nonlinear Hall resistances at the interface of Pt/FeSi bilayer, Phys. Rev. Mater. 8, 044407 (2024).
- Manyala et al. [2009] N. Manyala, B. D. Ngom, A. C. Beye, R. Bucher, M. Maaza, A. Strydom, A. Forbes, J. Johnson, A. T. Charlie, and J. F. DiTusa, Structural and magnetic properties of -Fe1-xCoxSi thin films deposited via pulsed laser deposition, Applied Physics Letters 94, 232503 (2009).
@@
- Breindel et al. [2023] A. Breindel, Y. Deng, C. M. Moir, Y. Fang, S. Ran, H. Lou, S. Li, Q. Zeng, L. Shu, C. T. Wolowiec, I. K. Schuller, P. F. S. Rosa, Z. Fisk, J. Singleton, and M. B. Maple, Probing FeSi, a d-electron topological Kondo insulator candidate, with magnetic field, pressure, and microwaves, Proceedings of the National Academy of Sciences 120, e2216367120 (2023).
@@
- Deng et al. [2023] Y. Deng, Y. Yan, H. Wang, E. Lee-Wong, C. M. Moir, Y. Fang, W. Xie, and M. B. Maple, Possible surface magnetism in the topological Kondo insulator candidate FeSi, Phys. Rev. B 108, 115158 (2023).
@@
- Brando et al. [2016] M. Brando, D. Belitz, F. M. Grosche, and T. R. Kirkpatrick, Metallic quantum ferromagnets, Rev. Mod. Phys. 88, 025006 (2016).
- Saha et al. [2022] S. Saha, A. Dutta, S. Gupta, S. Bandyopadhyay, and I. Das, Origin of the Griffiths phase and correlation with the magnetic phase transition in the nanocrystalline manganite , Phys. Rev. B 105, 214407 (2022).
- Das et al. [2018] K. Das, N. Banu, I. Das, and B. Dev, Magnetocaloric effect of polycrystalline Sm0.5Ca0.5MnO3 compound: Investigation of low temperature magnetic state, Solid State Communications 274, 36 (2018).
@@
- Ohnuma et al. [2012] I. Ohnuma, S. Abe, S. Shimenouchi, T. Omori, R. Kainuma, and K. Ishida, Experimental and Thermodynamic Studies of the Fe and Si Binary System, ISIJ International 52, 540 (2012).
@@
- Alexandrov et al. [2015] V. Alexandrov, P. Coleman, and O. Erten, Kondo Breakdown in Topological Kondo Insulators, Phys. Rev. Lett. 114, 177202 (2015).
@@
- Nakajima et al. [2016] Y. Nakajima, P. Syers, X. Wang, R. Wang, and J. Paglione, One-dimensional edge state transport in a topological Kondo insulator, Nature Physics 12, 213 (2016).
@@
- Sun et al. [2017] Y. Sun, Z. Zhuo, X. Wu, and J. Yang, Room-Temperature Ferromagnetism in Two-Dimensional Fe2Si Nanosheet with Enhanced Spin-Polarization Ratio, Nano Letters 17, 2771 (2017), pMID: 28441496, https://doi.org/10.1021/acs.nanolett.6b04884 .